A gas, initially at 292K, enclosed in a bulb of volume 15.1mL at a pressure of 90.2kPa, was first expanded into a reaction vessel so that the total volume was 182.2mL and then heated to 428K. What is the final pressure in the vessel?
1
Expert's answer
2020-09-29T06:12:16-0400

Solution.

"p1 \times V1 = p2\times V2"

"p2 = \frac{p1 \times V1}{V2}"

p2 = 7475.41 Pa

"\frac{p2}{T1} = \frac{p3}{T2}"

"p3 = \frac{p2 \times T2}{T1}"

p3 = 10957.11 Pa

Answer:

p3 = 10957.11 Pa